OpenStudy (samsan9):

If you save one penny on January 1, two pennies on January 2, three pennies on January 3, and continue this pattern for one year (not a leap year), what will be the value of your entire savings, in dollars, at the end of that one year? Express your answer as a decimal.

OpenStudy (anonymous):

$667.95?

OpenStudy (samsan9):

how did you get that?

OpenStudy (anonymous):

S = 1 + 2 + 3 + 4 +.....................................364 + 365 S =365+ 364+ 363+362+.........................................2 + 1 Adding 2S = 366 + 366 + 366 +...................365 times 2S = 365*366 S = 365*366/2

OpenStudy (anonymous):

it's mostly just a bunch of adding.

OpenStudy (samsan9):

did you add all that :o?

OpenStudy (anonymous):

no, there was a question just like this a while back.

OpenStudy (samsan9):

was it today?

OpenStudy (samsan9):

well thank you :D

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

yas!